OpenStudy (khally92):

A study has determined that the interaction at a large social party follow the mathematical progression N(t)=30t-t^2, where t is the time in minutes since the party began and N is the number of separate conversations occurring. at what time in a party do the most conversations occur? what is the maximum number of interactions? optimization problem, calculus.

OpenStudy (anonymous):

no, not calculus

OpenStudy (khally92):

optimization problem

OpenStudy (anonymous):

\[N(t)=30t-t^2\] is a parabola that opens down it will have max at the second coordinate of the vertex. first coordinate of the vertex of a parabola is \(-\frac{b}{2a}\) in your case \(-\frac{30}{2\times -1}=15\)

OpenStudy (anonymous):

so max occurs when \(t=15\) and the maximum number of interactions is \[N(15)=30\times 15-15^2\]

OpenStudy (anonymous):

you can use calc if you like, seems like overkill you would write \[N(t)=30t-t^2\] \[N'(t)=30-2t\] set is equal to zero to get the critical point write \[30-2t=0\]\[30=2t\]\[t=15\]

OpenStudy (anonymous):

in fact calc is a quick way to show that the first coordinate of the vertex of \(y=ax^2+bx+c\) is \(-\frac{b}{2a}\) \[y'=2ax+b\] set equal zero and solve, get the critical point at \(x=-\frac{b}{2a}\)

OpenStudy (khally92):

so t is 15 and N IS 225.